The boiling point of water is 100.00 °C at 1 atmosphere.

A student dissolves 13.51 grams of sodium sulfate, Na2SO4 (142.1 g/mol), in 262.4 grams of water. Use the table of boiling and freezing point constants to answer the questions below.


Solvent Formula Kb (°C/m) Kf (°C/m)
Water H2O 0.512 1.86
Ethanol CH3CH2OH 1.22 1.99
Chloroform CHCl3 3.67
Benzene C6H6 2.53 5.12
Diethyl ether CH3CH2OCH2CH3 2.02

The molality of the solution is _______ m.

The boiling point of the solution is __________ °C.
Answer #1

1)

Lets calculate molality first

mass(solute)= 13.51 g

use:

number of mol of solute,

n = mass of solute/molar mass of solute

=(13.51 g)/(1.421*10^2 g/mol)

= 9.507*10^-2 mol

m(solvent)= 262.4 g

= 0.2624 kg

use:

Molality,

m = number of mol / mass of solvent in Kg

=(9.507*10^-2 mol)/(0.2624 Kg)

= 0.3623 molal

Answer: 0.3623 molal

2)

i for Na2SO4 is 3 as 1 molecule of Na2SO4 dissociates into 2 Na+ and 1 SO42- ions

lets now calculate ΔTb

ΔTb = i*Kb*m

= 3.0*0.512*0.3623

= 0.5565 oC

This is increase in boiling point

boiling point of pure liquid = 100.0 oC

So, new boiling point = 100 + 0.5565

= 100.56 oC

Answer: 100.56 oC